OpenStudy (anonymous):

for a 14 inch pizza, find the area of a slice with 6/7 radians.

OpenStudy (anonymous):

Find the area of the pizza, then multiply by 6/7 divided by 2 pi, which is the portion of the disk in the piece.

OpenStudy (anonymous):

\[A_{slice}=49 \pi (\frac{6/7}{2\pi})=21 inches^2\]